OpenStudy (anonymous):

Which operation should be performed first according to the order of operations? 75 – 12 + 36 ÷ 3 + 14 • 8 A. 12 + 36 B. 75 – 12 C. 14 • 8 D. 36 ÷ 3

OpenStudy (anonymous):

@flower2468

OpenStudy (anonymous):

@mathstudent55

OpenStudy (k12awesomeness):

D.

OpenStudy (mathstudent55):

PEMDAS P - parentheses E - exponents MD - multiplications & divisions in the order they appear from left to right AS - additions and subtractions in the order they appear from left to right

OpenStudy (anonymous):

it's d

OpenStudy (k12awesomeness):

yep!

OpenStudy (mathstudent55):

P - there are no parentheses here E - there are no exponents here All we have is multiplications, divisions, additions, and subtractions.

OpenStudy (k12awesomeness):

yes

OpenStudy (mathstudent55):

Now start with MD - we do see a division written before a multiplication, so the first operation is the division, \(36 \div 3\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which operation should be performed first according to the order of operations? 44 • 2 – [(5 + 20 • 3) – 12] + 16 ÷ 2 A. 44 • 2 B. 5 + 20 C. 20 • 3 D. 16 ÷ 2

OpenStudy (anonymous):

i think it's b

OpenStudy (mathstudent55):

Here you have parentheses, so you need to work inside the parentheses first.

OpenStudy (texaschic101):

not b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c

OpenStudy (mathstudent55):

You start at the innermost parentheses. The curved parentheses inside the square brackets are the innermost parentheses.

OpenStudy (mathstudent55):

That means you need to start working on 5 + 20 • 3

OpenStudy (mathstudent55):

Now you look at that expression, and you apply PEMDAS to it. There are no longer any parentheses. There are no exponents, so you have MDAS left.

OpenStudy (mathstudent55):

MDAS means MD in order AS in order

OpenStudy (anonymous):

then ts d

OpenStudy (mathstudent55):

5 + 20 • 3 The 20 • 3 part is a multiplication, so that is done before the additions, so the answer is C.

OpenStudy (k12awesomeness):

Yes @pink33 it's C.

OpenStudy (mathstudent55):

The answer is C because you need to work in the parentheses first. Within the parentheses, you do the multiplication beofore the addition.

OpenStudy (anonymous):

Evaluate. 102 ÷ 25 • 20 + 4 ÷ 2 – 14

OpenStudy (anonymous):

1 more left

OpenStudy (mathstudent55):

44 • 2 – [(5 + 20 • 3) – 12] + 16 ÷ 2 = 44 • 2 – [(5 + 60) – 12] + 16 ÷ 2 = 44 • 2 – [65 – 12] + 16 ÷ 2 = 44 • 2 – 53 + 16 ÷ 2 = 88 – 53 + 16 ÷ 2 = 88 – 53 + 8 = 35 + 8 = 43

OpenStudy (anonymous):

so he answer is 43?

OpenStudy (k12awesomeness):

no, i'm guessing thats an example.

OpenStudy (anonymous):

well i guess i'll just guess the answer

OpenStudy (mathstudent55):

102 ÷ 25 • 20 + 4 ÷ 2 – 14 Do you see parentheses or exponents?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

28.8 it is

OpenStudy (k12awesomeness):

yep

OpenStudy (texaschic101):

weird one...im ending up with a decimal

OpenStudy (mathstudent55):

Ok, so we don't have PE in PEMDAS. Now we move on to the next level. MD Do you see multiplications and additions?

OpenStudy (mathstudent55):

There is nothing wrong with decimals, and in fact the answer here is a decimal.

OpenStudy (k12awesomeness):

yeah, @pink33 already knows it's 28.8

OpenStudy (texaschic101):

but not 28.8

OpenStudy (mathstudent55):

\(102 \div 25 \cdot 20 + 4 \div 2 - 14\) It is not 28.8

OpenStudy (k12awesomeness):

well, what did u get @texaschic101 ?

OpenStudy (mathstudent55):

We are working on the MD level of PEMDAS. What is the first multiplication or division you see from left to right?

OpenStudy (texaschic101):

69.6

OpenStudy (anonymous):

thns i failed yaay!

OpenStudy (k12awesomeness):

i don't see how you got 69.6 and I also looked it up online and I'm still getting 28.8 or if you round 29

OpenStudy (mathstudent55):

We can do it together. \(102 \div 25 \cdot 20 + 4 \div 2 - 14\)

OpenStudy (mathstudent55):

What is the first operation you need to do?

OpenStudy (texaschic101):

102/25 = 4.08 4.08 * 20 = 81.6 + 2 - 14 = 69.6

OpenStudy (mathstudent55):

P - no parentheses here E - no exponents here MD - we do have multiplications and divisions What is the first division?

OpenStudy (k12awesomeness):

ohh, i see now.

OpenStudy (k12awesomeness):

i hit a wrong key in my calculator.

OpenStudy (mathstudent55):

\(102 \color{red}{\div} 25 \cdot 20 + 4 \div 2 - 14\) \(= 4.08 \color{red}{\cdot} 20 + 4 \div 2 - 14\) \(= 81.6 + 4 \color{red}{\div} 2 - 14\) \(= 81.6 \color{red}{+} 2 - 14\) \(= 83.6 \color{red}{-} 14\) \(= 69.6 \)